OpenStudy (anonymous):

There are three different positive integers whose reciprocals add up to 1. What is the product of those three integers?

OpenStudy (anonymous):

\[\frac{1}{x}+\frac{1}{y}+\frac{1}{z}=1\] where \(x,y,z\) are the three positive integers. \[\begin{align*}\frac{yz}{xyz}+\frac{xz}{xyz}+\frac{xy}{xyz}&=1\\\\ yz+xz+xy&=xyz \end{align*}\] The first solution that comes to mind is \(x=2\), \(y=3\) and \(z=6\), which gives \(xyz=36\). Not sure if there are more possibilities for \(x,y,z\)...

OpenStudy (anonymous):

OH MY GOSH thanks so much it's right your my hero!!!!!!!!!!!!!!!

OpenStudy (anonymous):

thanks:)

OpenStudy (anonymous):

yw